# Taming Uncertainty via Automation: Observing, Analyzing, and Optimizing Agentic AI Systems

### v1

@@ Abstract
This paper introduces AgentOps: a comprehensive framework for observing, analyzing, optimizing, and automating operation of agentic AI systems. We identify distinct needs across four key roles—developers, testers, site reliability engineers (SREs), and business users—each of whom engages with the system at different points in its lifecycle. We present the AgentOps Automation Pipeline, a six-stage process encompassing behavior observation, metric collection, issue detection, root cause analysis, optimized recommendations, and runtime automation. Throughout, we emphasize the critical role of automation in managing uncertainty and enabling self-improving AI systems—not by eliminating uncertainty, but by taming it to ensure safe, adaptive, and effective operation.
Index Terms:
Large Language Models, Multi-Agent Systems, Monitoring, Analytics, Observability, Agentic systems, Performance Optimization, Evaluation
@@ I Introduction
Designed for autonomy, agentic systems move beyond fixed instructions to make context-aware decisions using probabilistic reasoning. Their behavior is shaped by LLMs, classical AI planning, and machine learning, often producing divergent outcomes for identical inputs [1].
@@
This dynamic landscape introduces significant challenges for users—developers, evaluators, SREs, and business stakeholders—across the system lifecycle. Surveys show that only 8% of organizations use dedicated observability platforms [2], limiting automation and scalability, and 60% of users report that current analytics tools do not meet their needs [3].
Addressing this uncertainty demands more than traditional operations. It requires a new discipline: AgentOps [4, 5, 6].
@@ II Related Work: Landscape of Agent Analytics
The agent observability and analytics landscape has split into two categories: GenAI tools like Phoenix [7], LangFuse [8], and LangSmith [9] targeting developers, and observability platforms like Datadog [10] and IBM Instana [11], which are evolving to support agentic systems for SREs.
All of these rely on observability and data collection, reinforcing the need for standardized protocols. OpenTelemetry (OTel)[12], a key standard for logs, traces, and metrics, is being extended to support agent-based workflows. OpenLLMetry[13] by Traceloop enables observability for frameworks like LangGraph [14], CrewAI [15], and AutoGen [16]. Other notable efforts include OpenInference[17] and LangFuse[8]. Still, there are no widely adopted semantic conventions for agentic tracing, and instrumentation for behaviors like planning or reflection remains limited.
Beyond observability, analytics tools lack standardized failure taxonomies. While early proposals exist [18, 19], they are not widely adopted. Critical issues—like unintended loops in multi-agent workflows—often go undetected. Graph Neural Networks [20] offer a promising direction for encoding agentic structure, but failure analysis using them is still nascent.
Root cause analysis remains similarly underdeveloped. Few existing tools effectively capture causal relationships between agent decisions, tool behaviors, and observed failures. Prior work on causal reasoning in process mining [21] and causal discovery in agentic systems [22] offers promising foundations for adaptation in this context.
@@
Finally, automation remains limited. Most systems require manual intervention, even for recurring issues. While monitoring exists, agentic workflows are rarely refined or adapted automatically.
@@ III-B Testers
III-C Software Reliability Engineers
@@ IV-F Automate Operations
For example, AgentOps observes behavior by noticing that an SRE Agent [23] inconsistently uses a new diagnostic tool, based on traces of its workflow execution and tool invocations. AgentOps collects metrics, monitoring issue rates and identifying frequent incomplete outputs for a specific task. Next, AgentOps detects issues by spotting a drift in this failure pattern, indicating rising task failures. AgentOps identifies the root cause as inaccurate prompt instructions and tool misuse. It then optimizes recommendations by suggesting a clearer prompt. Finally, AgentOps automates operation by augmenting the Agent’s prompt and revalidating the fix through continuous monitoring—automatically optimizing behavior without changing code. Like a manager guiding an employee based on performance feedback, AgentOps enables agentic systems to self-correct and improve in real time.
@@ V Discussion: Taming, not Eliminating, Uncertainty
  • Standardization. Our taxonomy provides a foundation for AgentOps instrumentation, evaluation, and automation, aligned with emerging standards such as OpenTelemetry, MCP [24], and the UIM Protocol [25] that promote interoperability across agentic systems.

### v2

@@ Abstract
This paper presents our vision of AgentOps: a comprehensive framework for observing, analyzing, optimizing, and automating operation of agentic AI systems. We identify distinct needs across four key roles—developers, testers, site reliability engineers (SREs), and business users—each of whom engages with the system at different points in its lifecycle. We present the AgentOps Automation Pipeline, a six-stage process encompassing behavior observation, metric collection, issue detection, root cause analysis, optimized recommendations, and runtime automation. Throughout, we emphasize the critical role of automation in managing uncertainty and enabling self-improving AI systems—not by eliminating uncertainty, but by taming it to ensure safe, adaptive, and effective operation.
footnotetext: International Conference on Automated Software Engineering, ASE 2025.
@@ I Introduction
Designed for autonomy, agentic systems move beyond fixed instructions to make context-aware decisions using probabilistic reasoning. Their behavior is shaped by LLMs, classical AI planning, and machine learning, often producing divergent outcomes for identical inputs [atil2024non].
@@
This dynamic landscape introduces significant challenges for users—developers, evaluators, SREs, and business stakeholders—across the system lifecycle. Surveys show that only 8% of organizations use dedicated observability platforms [precisely, precisely_report], limiting automation and scalability, and 60% of users report that current analytics tools do not meet their needs [moshkovich2025beyond].
Addressing this uncertainty demands more than traditional operations. It requires a new discipline: AgentOps [agentops_ai, ibm_agentops, dong2024taxonomy].
@@ II Related Work: Landscape of Agent Analytics
The agent observability and analytics landscape has split into two categories: GenAI tools [jose2024harnessing] like Phoenix [arize], LangFuse [langfuse], and LangSmith [langsmith] targeting developers, and observability platforms like Datadog [datadog] and IBM Instana [instana], which are evolving to support agentic systems for SREs.
All of these rely on observability and data collection, reinforcing the need for standardized protocols. OpenTelemetry (OTel)[blanco2023practical], a key standard for logs, traces, and metrics, is being extended to support agent-based workflows. OpenLLMetry[openllmetry] by Traceloop enables observability for frameworks like LangGraph [langgraph], CrewAI [crewai], and AutoGen [autogen]. Other notable efforts include OpenInference[openinference] and LangFuse[langfuse]. Still, there are no widely adopted semantic conventions for agentic tracing, and instrumentation for behaviors like planning or reflection remains limited [moshkovich2025otel].
Beyond observability, analytics tools lack standardized failure taxonomies. While early proposals exist [deshpande2025trail, pan2025multiagent], they are not widely adopted. Critical issues—like unintended loops in multi-agent workflows—often go undetected. Graph Neural Networks [niro2023detecting] offer a promising direction for encoding agentic structure, but failure analysis using them is still nascent.
Root cause analysis remains similarly underdeveloped. Few existing tools effectively capture causal relationships between agent decisions, tool behaviors, and observed failures. Prior work on causal reasoning in process mining [fournier2025business] and causal discovery in agentic systems [fournier2025agentic] offers promising foundations for adaptation in this context.
@@
Finally, automation remains limited, with only initial efforts to design approaches tailored to agentic systems [sanwouo2025breaking]. Most systems require manual intervention, even for recurring issues, and workflows are rarely refined or adapted automatically.
@@ III-B Testers
III-C Site Reliability Engineers
@@ IV-F Automate Operations
For example, AgentOps observes behavior by noticing that an SRE Agent [jha2025itbench] inconsistently uses a new diagnostic tool, based on traces of its workflow execution and tool invocations. AgentOps collects metrics, monitoring issue rates and identifying frequent incomplete outputs for a specific task. Next, AgentOps detects issues by spotting a drift in this failure pattern, indicating rising task failures. AgentOps identifies the root cause as inaccurate prompt instructions and tool misuse. It then optimizes recommendations by suggesting a clearer prompt. Finally, AgentOps automates operation by augmenting the Agent’s prompt and revalidating the fix through continuous monitoring—automatically optimizing behavior without changing code. Like a manager guiding an employee based on performance feedback, AgentOps enables agentic systems to self-correct and improve in real time.
@@ V Discussion: Taming, not Eliminating, Uncertainty
  • Standardization. Our taxonomy provides a foundation for AgentOps instrumentation, evaluation, and automation, aligned with emerging standards such as OpenTelemetry, MCP [hou2025mcp], and the UIM Protocol [uim] that promote interoperability across agentic systems.